Santa's fish tank contains 200 gallons of sea water. The sea water is 7.5% salt. How many gallons of fresh water must be added to the tank so that the mixture only contains 7% salt?
Please help! I'm bad at mixtures and setting them up.

Answer by stanbon(75887)

Santa's fish tank contains 200 gallons of sea water. The sea water is 7.5% salt. How many gallons of fresh water must be added to the tank so that the mixture only contains 7% salt?
-------
Equation:
salt + salt = salt
----
0.075*200 + 0*x = 0.07(200+x)
-----
Multiply thru by 1000 to get:
75*200 + 0 = 70*200 + 70x
---
70x = 5*200
x = 14.29 gallons (amount of water that must be added)
===============
Cheers,
Stan H.
===============